# Electromagnetic Flow Control in Hypersonic Rarefied Environment

### v1

@@ 4 Conclusions
In this study, the UGKWP method has been employed for the multiscale simulation of partially ionized plasma flows around a hemisphere, spanning regimes from near-continuum to highly rarefied. Validation against established UGKS codes for neutral flows, as well as experimental data for pre-ionized argon flows, demonstrates the accuracy and reliability of our numerical implementation. The results reveal that rarefied effects have a substantial impact on the prediction of electromagnetic flow control, underscoring the necessity of multiscale modeling in plasma flow applications.
Future work will focus on incorporating additional physical phenomena, such as ambipolar diffusion, shock polarization, wall sheath effects arising from small Debye lengths, and comprehensive magnetohydrodynamic effects for moderate to large magnetic Reynolds numbers. The fidelity of these advanced simulations will continue to depend on the robust flow field solutions provided by the UGKWP multiscale solver. Overall, this research establishes a solid foundation for the development of more sophisticated plasma flow simulations in the future.
@@ 5 Acknowledgement
This current work was supported by National Key R&\&&D Program of China (Grant Nos. 2022YFA1004500), National Natural Science Foundation of China (12172316, 92371107), and Hong Kong research grant council (16301222, 16208324).

### v2

@@ 4 Conclusions
This study applies the UGKWP method to multiscale simulations of partially ionized plasma flows around a hemisphere, spanning regimes from near-continuum to highly rarefied. Validation against established UGKS solutions for neutral flows, along with experimental data for pre-ionized argon flows, confirms the accuracy and robustness of the numerical implementation. A comparative investigation of control effects across different Knudsen numbers shows that rarefied effects can significantly alter predictions of electromagnetic flow control. Future work will extend the framework to include additional physical phenomena, such as wall sheath dynamics and comprehensive magnetohydrodynamic effects at moderate to high magnetic Reynolds numbers. The reliability of these advanced simulations will continue to rely on the robust multiscale flow-field solutions enabled by the UGKWP solver. Taken together, the present study establishes a solid foundation for advancing the simulation of complex plasma flow phenomena.
@@ 5 Acknowledgement
This current work was supported by National Key R&\&D Program of China (Grant Nos. 2022YFA1004500), National Natural Science Foundation of China (12172316, 92371107), and the Hong Kong Research Grant Council (16301222, 16208324).
Appendix A Artificial electron mass effect
According to the Drude’s conductivity model [33],
$\sigma^{cond}=\frac{n_{e}e^{2}}{m_{e}\sum_{s\neq e}\nu_{e,s}},$
where $\nu_{e,s}$ is the collision frequency between species $s$ and electrons and is given as [34],
$\nu_{e,s}=n_{s}\sigma_{e,s}^{col}\bar{v}_{th}$
where $n_{s}$ is number density of species $s$, $\sigma_{e,s}^{col}$ is the cross section between species $s$ and electrons. $\bar{v}_{th}$ is the average relative velocity between species $s$ and electrons and is given approximately as the average velocities of electrons $\bar{v}_{th}=\sqrt{\frac{8k_{B}T_{e}}{\pi m_{e}}}$. Therefore, the conductivity is given as
$\sigma^{cond}=\frac{n_{e}e^{2}}{m_{e}\sum_{s\neq e}n_{s}\sigma_{e,s}^{col}\sqrt{\frac{8k_{B}T_{e}}{\pi m_{e}}}}.$ (15)
In the current work, $\sigma_{e,s}^{col}$ is given by a fitted curve, and $n_{e},T_{e}$ and $n_{s}$ are calculated by the flow field so their expressions don’t contain $m_{e}$, then we have
$\sigma^{cond}\propto\frac{1}{\sqrt{m_{e}}},$
which means conductivity decreases as $m_{e}$ increases. Standoff distance is proportional to the dimensionless magnetic interaction parameter [7],
$Q=\frac{\sigma^{cond}_{0}|\boldsymbol{B}_{0}|^{2}L_{0}}{\rho_{0}|\boldsymbol{U}|_{0}},$ (16)
where subscript 0 means characteristic quantities. To make $Q$ constant, $|\boldsymbol{B}_{0}|^{2}$ has to increase when the conductivity decreases. Given that the actual mass ratio of argon ions ($Ar^{+}$) to electrons ($e^{-}$) is approximately 72811, while the ratio employed in this paper is 10, the electrical conductivity is consequently reduced by a factor of approximately 85, $|\boldsymbol{B}_{0}|^{2}$ needs to be increased by a factor of 85 to maintain the desired similitude.
Appendix B Discretizations of cross-species collision term
The discretized form of Eq.14 is
$\frac{\boldsymbol{U}_{\alpha}^{n+1}-\boldsymbol{U}_{\alpha}^{**}}{\Delta t}=\sum_{k=1}^{N}\frac{2m_{k}\chi_{\alpha k}n_{k}}{m_{\alpha}+m_{k}}(\boldsymbol{U}_{k}^{n+1}-\boldsymbol{U}_{\alpha}^{n+1}),$
$\frac{E_{\alpha}^{n+1}-E_{\alpha}^{**}}{\Delta t}=\sum_{k=1}^{N}\frac{2m_{k}\chi_{\alpha k}n_{k}}{m_{\alpha}+m_{k}}\boldsymbol{U}_{\alpha}^{n+1}\left(\boldsymbol{U}_{k}^{n+1}-\boldsymbol{U}_{\alpha}^{n+1}\right)$
$\;+\sum_{k=1}^{N}\frac{4m_{k}\chi_{\alpha k}n_{k}}{(m_{\alpha}+m_{k})^{2}}\left[m_{k}(E_{k}^{n+1}-\frac{1}{2}(\boldsymbol{U}_{k}^{n+1})^{2})-m_{\alpha}(E_{\alpha}^{n+1}-\frac{1}{2}(\boldsymbol{U}_{\alpha}^{n+1})^{2})+\frac{m_{k}}{2}(\boldsymbol{U}_{k}^{n+1}-\boldsymbol{U}_{\alpha}^{n+1})^{2}\right].$
The Jacobian of residuals to velocity and energy is
$\frac{\partial R_{U_{\alpha}}}{\partial\boldsymbol{U}_{k}}=\delta_{\alpha k}(\frac{1}{\Delta t}+\sum_{k=1}^{N}\frac{2m_{k}\chi_{\alpha k}n_{k}}{m_{\alpha}+m_{k}})-(1-\delta_{\alpha k})(\frac{2m_{k}\chi_{\alpha k}n_{k}}{m_{\alpha}+m_{k}}),$
$\frac{\partial R_{E_{\alpha}}}{\partial E_{k}}=\delta_{\alpha k}(\frac{1}{\Delta t}+\sum_{k=1}^{N}\frac{4m_{k}m_{\alpha}\chi_{\alpha k}n_{k}}{(m_{\alpha}+m_{k})^{2}})-(1-\delta_{\alpha k})\frac{4m_{k}^{2}\chi_{\alpha k}n_{k}}{(m_{\alpha}+m_{k})^{2}},$
$\frac{\partial R_{E_{\alpha}}}{\partial\boldsymbol{U}_{k}}=\delta_{\alpha k}\sum_{k=1}^{N}\frac{4m_{k}\chi_{\alpha k}n_{k}}{(m_{\alpha}+m_{k})^{2}}\left(-m_{\alpha}\boldsymbol{U}_{\alpha}+m_{k}(\boldsymbol{U}_{k}-\boldsymbol{U}_{\alpha})\right)+$
$\quad(1-\delta_{\alpha k})(\frac{4m_{k}^{2}\chi_{\alpha k}n_{k}}{(m_{\alpha}+m_{k})^{2}}\left(\boldsymbol{U}_{k}-m_{k}(\boldsymbol{U}_{k}-\boldsymbol{U}_{\alpha})\right))$
where
$\delta_{\alpha k}=\left\{\begin{array}[]{cc}1&\text{if}\alpha=k\\ 0&\text{else}\end{array}\right$.
Residuals are
$R_{U_{\alpha}}=\frac{\boldsymbol{U}_{\alpha}^{n+1}-\boldsymbol{U}_{\alpha}^{**}}{\Delta t}-\sum_{k=1}^{N}\frac{2m_{k}\chi_{\alpha k}n_{k}}{m_{\alpha}+m_{k}}(\boldsymbol{U}_{k}^{n+1}-\boldsymbol{U}_{\alpha}^{n+1}),$
$R_{E_{\alpha}}=\frac{E_{\alpha}^{n+1}-E_{\alpha}^{**}}{\Delta t}-\sum_{k=1}^{N}\frac{2m_{k}\chi_{\alpha k}n_{k}}{m_{\alpha}+m_{k}}\boldsymbol{U}_{\alpha}^{n+1}\left(\boldsymbol{U}_{k}^{n+1}-\boldsymbol{U}_{\alpha}^{n+1}\right)$
$\;+\sum_{k=1}^{N}\frac{4m_{k}\chi_{\alpha k}n_{k}}{(m_{\alpha}+m_{k})^{2}}\left[m_{k}(E_{k}^{n+1}-\frac{1}{2}(\boldsymbol{U}_{k}^{n+1})^{2})-m_{\alpha}(E_{\alpha}^{n+1}-\frac{1}{2}(\boldsymbol{U}_{\alpha}^{n+1})^{2})+\frac{m_{k}}{2}(\boldsymbol{U}_{k}^{n+1}-\boldsymbol{U}_{\alpha}^{n+1})^{2}\right].$
Then iterative solvers like the Newton method can be used to solve the nonlinear system.